I need to write a system of "two" equations that have:
1. a unique solution
2. an infinate solution
3. no solution

Answer by Edwin McCravy(20055)


The system of two equations

2x - 3y = 13
5x + 2y =  4

have a unique solution (x,y) = (2,-3)

The system of two equations:

 4x - 3y =  3
 6y - 8x = -6 

Has infinitely many solutions.  For example

(x,y) = (-6,-9), (-3,-5), (0,-1), (3,3, (6,7) and
millions and billions more solutions


The syatem of two equations:

  3x + 7y = 9
-14y - 6x = 8

has no solution  

Edwin